# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ 2.2 Separation between inductive and transductive settings
Proposition 2.1.
Let $d=1$. Consider the algorithm, which computes, upon observing $x_{t}$ at time $t$,
@@
Proof of Proposition 2.1 We consider two cases. In the first case, the regularization is given at a large scale $\lambda\geq T$. For this case, consider $x_{t}=1$ and $y_{t}=1$ for all $t$. Then we have
@@
$=\left\lvert 1+\frac{T^{2}2^{t+1}\left(2^{t-1}+(-1)^{t}\right)/3}{\lambda+4T^{2}(4^{t}-1)/3}\right\rvert$
@@
Although our result does not resolve the question in (Gaillard et al., 2019, Section 4.1), it hints, in view of Theorem 2.1, at a separation between the transductive setup and the inductive setup. In the following section, we show that the separation indeed exists for the clipped linear function class.
@@
$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\inf_{\theta\in\mathbb{R}}\left(\sum_{t=1}^{T}(y_{t}-\operatorname{clip_{m}}\left(x_{t}\theta\right))^{2}\right)\geq\Omega(T).$
Proof of Proposition 2.2 With a slight abuse of the notation, denote $x^{(i)}=10^{i}$ and $\theta^{(i)}=\frac{15}{16x^{(i)}}$ for $i\in\{0,1,\ldots,2^{T}-2\}$. We have
$\begin{cases}\theta^{(i)}x^{(j)}>7/8&\text{if~{}}j\geq i,\\ \theta^{(i)}x^{(j)}<1/8&\text{if~{}}j<i.\end{cases}$
@@
$\begin{split}\begin{cases}x_{t}\widetilde{\theta}>y_{t}=7/8&\text{if~{}}\hat{y}_{t}\leq 1/2,\\ x_{t}\widetilde{\theta}<y_{t}=1/8&\text{if~{}}\hat{y}_{t}>1/2\end{cases}\end{split}$ (17)
with the following binary searching construction. To start with, we set the left and the right end to be respectively $a_{1}=0$ and $b_{1}=2^{T}-2$. At each round $t$, the adversary presents the middle point of $a_{t}$ and $b_{t}$, that is, $i_{t}=\lfloor\frac{a_{t}+b_{t}}{2}\rfloor$ and $x_{t}=x^{(i_{t})}$. Then if $\hat{y}_{t}\leq 1/2$, then we set $y_{t}=7/8$, $a_{t+1}=x^{(i_{t}+1)}$, and $b_{t+1}=b_{t}$, else we set $y_{t}=1/8$, $a_{t+1}=a_{t}$, and $b_{t+1}=x^{(i_{t}-1)}$. At the beginning, we have $n_{1}=2^{T}-1$ covariates in between $a_{1}$ and $b_{1}$ and each time we are left with $(n_{t}-1)/2$ covariates in between $a_{t+1}$ and $b_{t+1}$. Then by induction, $n_{t}=2^{T-t+1}-1$. Thus, $n_{T}=1$. Finally, it is clear that $\widetilde{\theta}=\theta^{(i_{T})}$ satisfies (17).
@@ 3 Logistic Regression
[…]
@@
We provide the first non-asymptotic, tight upper bound for the minimax excess risk for transductive logistic regression without any assumptions on the covariates $x_{1},\dots,x_{T}$ or the optimal comparator ${\theta^{\star}}$. Importantly, in Section 5, we discuss implementable versions of our algorithms that are guaranteed to converge to the optimal solution, with the only caveat that the number of iterations required may grow exponentially with the dimension in the worst-case scenario. In the specific case of random design logistic regression, we achieve in Section 6.2 a truly polynomial time implementation.
@@ 3.1 Background for exponential weights algorithm for logistic loss
Similar to the mixability of the squared loss (Lemma 2.3), the mixability of logistic loss facilitates the analysis of exponential weights algorithms (Vovk, 1990), or equivalently, Bayesian learning algorithms (Kakade and Ng, 2004). Formally, the logistic loss defined as $\ell_{\theta}(x,y)=-\log(\sigma(y\langle x,\theta\rangle))$ is $1$-mixable, i.e., for any $\rho$, $x$, and $y$,
$-\log\left(\mathbb{E}_{\theta\sim\rho}\left[\sigma(y\langle x,\theta\rangle)\right]\right)=-\log\left(\mathbb{E}_{\theta\sim\rho}\exp(-\eta\ell_{\theta}(x,y))\right).$ (18)
@@
[…]
where ${\theta^{\star}}\vcentcolon=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum\limits_{t=1}^{T}\ell_{\theta}(x_{t},y_{t})$ and $\mathsf{H}_{t}({\theta^{\star}})$ is the Hessian matrix of $\ell_{\theta}(x_{t},y_{t})$ at ${\theta^{\star}}$. Applying an exact Laplace formula for quadratic forms (Lemma 3.3 below) to (LABEL:eq:laplace-approximation) together with Lemma 2.1 and Eq. 18, the asymptotic logistic loss of an exponential weights algorithm with prior $\mu$ is approximately
@@
Instead, Kakade and Ng (2004) employs a Gaussian prior and obtains an upper bound through quadratic approximation of the logistic loss. Concretely, by choosing the prior distribution $\mu=\mathcal{N}(0,\nu^{2}I)$ to be gaussian for some $\nu>0$ to be specified, by Lemmas 2.1, 2.2 and 18, the loss of the exponential weights algorithm satisfies
$-\sum\limits_{t=1}^{T}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\left[\sigma(y_{t}\langle x_{t},\theta\rangle)\right]\right)=\inf_{\varphi}\left(-\sum\limits_{t=1}^{T}\mathbb{E}_{\theta\sim\varphi}\left[\sigma(y_{t}\langle x_{t},\theta\rangle)\right]+\mathcal{KL}(\varphi\|\mu)\right).$
By choosing $\varphi=\mathcal{N}({\theta^{\star}},\varepsilon^{2}I)$ for some $\varepsilon>0$ to be specified, together with $|\partial^{2}(-\log\sigma(y_{t}z))/\partial z^{2}|\leq 1/4$, the quadratic approximation of $-\mathbb{E}_{\theta\sim\varphi}\sum\limits_{t=1}^{T}\log\sigma(y_{t}\langle x_{t},\theta\rangle)$ near ${\theta^{\star}}$ is upper bounded by
@@
In fact, ${\theta^{\star}}$ can be unbounded in the logistic regression case. Moreover, even $\min_{t}|\langle{\theta^{\star}},x_{t}\rangle|$ can be unbounded, as can be seen from the realizable case where all $x_{t}$-s have the same label. Indeed, since a lower bound for the inductive case is established by Foster et al. (2018, Thereom 5) in the following theorem, it is not possible to obtain our result with any prior that does not depend on $x_{1},\ldots,x_{T}$.
Theorem 3.1 (Foster et al. (2018)).
@@
This theorem, in conjunction with our upper bound Theorem 3.2, establishes a learnability separation in terms of assumptions between the transductive and inductive settings for logistic loss. Specifically, it demonstrates that the transductive setting is crucial for achieving assumption-free regret bounds.
@@ 3.2 Slab-experts
For this comparator [^2] The parameter ${\theta^{\star}}$ is not always well-defined in $\mathbb{R}^{d}$. For instance, when the sequence is realizable, ${\theta^{\star}}$ will tend to be infinite in the separating direction. Since this is a minor technical issue, we avoid the complication in this paper. For a more rigorous treatment, we refer to Mourtada and Gaïffas (2022). and any $1\leq t\leq T$, if $y_{t}\langle x_{t},{\theta^{\star}}\rangle\leq 0$, then $|\langle{\theta^{\star}},x_{t}\rangle|\leq T$.
Proof of Lemma 3.1 For any $t$,
$-\log(\sigma(y_{t}\langle x_{t},{\theta^{\star}}\rangle))\leq\sum_{t=1}^{T}-\log(\sigma(y_{t}\langle x_{t},{\theta^{\star}}{}\rangle))\leq\sum_{t=1}^{T}-\log(\sigma(0))=T\log 2,$
@@
Lemma 3.1 states that for any $x_{t}$ wrongly predicted covariate by ${\theta^{\star}}$, the margin $|\langle{\theta^{\star}},x_{t}\rangle|$ will not exceed $T$. This motivates the following separation of the covariates: either $|\langle{\theta^{\star}},x_{t}\rangle|\leq T$ or $y_{t}\langle{\theta^{\star}},x_{t}\rangle>T$. In the following, we first introduce a partial exponential weights algorithm that depends on any chosen subsets of the covariates that only predicts on the chosen subsets. Then, we use these partial exponential weights algorithms to build slab-experts.
Definition 3.1 (Partial exponential weights).
@@
Definition 3.2 (Slab-expert).
A slab-expert $B_{\theta,\alpha}$ is an expert that is specified by a vector $\theta\in\mathbb{R}^{d}$ and a regularization parameter $\alpha>0$. Given $x_{1},\ldots,x_{T}$, the expert constructs a slab $K_{\theta}=\{x_{t}:\left\lvert\langle x_{t},\theta\rangle\right\rvert\leq T\}$ and predicts labels in the following way: At times step $t$, if $x_{t}\in K_{\theta}$, the expert predicts on $x_{t}$ according to $\mathrm{EWA}_{K_{\theta},\alpha}$ with logistic loss and learning rate $1$. If $x_{t}\notin K_{\theta}$, then the expert predicts $\hat{y}_{t}=\operatorname{sign}(\langle x_{t},\theta\rangle)$ with probability $1$. We denote by $B_{\theta,\alpha}(x_{t},y_{t})$ the probability of predicting $y_{t}$ upon observing $x_{t}$.
@@
Construct the set of all slab-experts $\mathcal{A}=\{B_{\theta,\alpha}\mid{}\theta\in\mathbb{R}^{d}\}$ with logistic loss and learning rate $1$.
@@
      The learner predicts with aggregation on the slab-experts $\tilde{\rho}_{t}(x_{t},\cdot)=\sum\limits_{B\in\mathcal{A}}p_{t}(B)B(x_{t},\cdot)\text{~{}~{}where~{}~{}}p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}-\log B(x_{i},y_{i})\right)$ (22) and $B(x_{i},y_{i})$ is probability of $B$ predicting $y_{i}$ on $x_{i}$.
@@
$-\sum\limits_{t=1}^{T}\log\tilde{\rho}_{t}(x_{t},y_{t})\leq-\sum\limits_{t=1}^{T}\log(\sigma(y_{t}\langle x_{t},\theta^{\star}\rangle))+11d\log(eT).$
There are several noteworthy aspects of this result. First and foremost, it provides the first assumption-free guarantee, meaning that no assumptions are required on the covariates $x_{1},\dots,x_{T}$ or the comparator ${\theta^{\star}}$. In Appendix D, we will present an alternative covering approach that achieves the same upper bound, where the covering is constructed in a manner analogous to the slab division on the covariates. However, our focus remains on the slab-experts approach, as the covering approach lacks implementable algorithms for constructing the covering. Indeed, our algorithm can be implemented in time polynomial in $T$, though exponential in $d$ in the worst case. This efficiency stems from the slab-experts themselves, each of which consists of an exponential weight algorithm with a Gaussian prior and a deterministic linear separator. The exponential weight algorithm with a Gaussian prior can be computed in polynomial time through log-concave sampling, and the aggregation of the slab-experts can be performed in $\mathrm{poly}(T^{d})$ time. Importantly, under mild additional assumptions on the distribution in the i.i.d. batch setup, we can reduce the number of experts to $2$ with pertaining, providing a polynomial time algorithm. For further details, see Sections 5 and 6. Additionally, the slab-experts approach has two notable aspects:
@@
Lemma 3.2.
There are at most $(eT/d)^{9d}$ distinct slab-experts.
Proof of Lemma 3.2 We note that by Blumer et al. (1989, Lemma 3.2.3), the VC dimension of the intersection of two half-spaces is at most $8d$. So the number of all possible subsets $K_{\theta}$ is upper bounded by $(eT/d)^{8d}$ by the Vapnik-Chervonenkis-Sauer-Shelah lemma (Shalev-Shwartz and Ben-David (2014, Lemma 6.10)). Let all the subsets be enumerated as $K_{1},K_{2},\dots,K_{J}$. For each subset $K_{j}$ for $j\in[J]$ and any vector $\theta$ such that $K_{\theta}=K_{j}$, since the slab-expert $K_{\theta}$ is realizable on points outside the slab, i.e., $K_{\theta}^{c}=\{x_{t}\}_{t\in[T]}\setminus K_{\theta}$. Again, by the Vapnik-Chervonenkis-Sauer-Shelah lemma, there are no more than $(eT/d)^{d}$ possibilities on the set $K_{\theta}^{c}$. Altogether, there are no more than $(eT/d)^{9d}$ slab-experts. ∎
@@
Proposition 3.1.
Let $K\subset\{x_{t}\}_{t\in[T]}$ and ${\theta^{\star}}\in\mathbb{R}^{d}$. Then the partial exponential weights algorithm $\mathrm{EWA}_{K,\alpha}$ as defined in Definition 3.1 with prediction restricted on covariates in $K$ satisfies a loss upper bound of
$-\sum_{x\in K}\log\left(\mathbb{E}_{\theta\sim\rho_{x}}\left[\sigma(y_{x}\langle x,\theta\rangle)\right]\right)\leq-\sum_{x\in K}\log(\sigma(y_{x}\langle x,\theta^{\star}\rangle))+\alpha({\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right){\theta^{\star}}+\frac{d}{2}\log\left(1+\frac{1}{8\alpha}\right)$
on the set $K$ where $\rho_{x}$ is the prediction of $\mathrm{EWA}_{K,\alpha}$ on $x$. Specifically, suppose that ${\theta^{\star}}$ as defined in (21) satisfies $|\langle x,{\theta^{\star}}\rangle|\leq T$ for all $x\in K$ and let $\alpha=1/T^{3}$. We have
@@
Lemma 3.3 (Exact Laplace formula for quadratic forms).
Let $Q(\theta)=\theta^{\top}A\theta+b^{\top}\theta+c$, where $A\in\mathbb{R}^{d\times d}$ is a positive semi-definite matrix, $b\in\mathbb{R}^{d}$ and $c\in\mathbb{R}$. Then
[…]
Proof of Proposition 3.1 By Lemma 2.1, we have
$-\sum_{x\in K}\log\left(\mathbb{E}_{\theta\sim\rho_{x}}\left[\sigma(y_{x}\langle x,\theta\rangle)\right]\right)=-\log\left(\int\limits_{\mathbb{R}^{d}}\exp\left(\sum\limits_{x\in K}\log(\sigma(y_{x}\langle x,\theta\rangle))\right)\mu_{K}(d\theta)\right).$
Since $|f^{\prime\prime}(z)|\leq 1/4$, using the Taylor expansion for $f$ we have for any $x\in K$,
$-\log(\sigma(y_{x}\langle x,\theta\rangle))\leq-\log(\sigma(y_{x}\langle x,{\theta^{\star}}\rangle))-\frac{1}{1+\exp(y_{x}\langle x,{\theta^{\star}}\rangle)}\cdot(y_{x}\langle x,\theta-{\theta^{\star}}\rangle)+\frac{1}{8}(\langle x,\theta-{\theta^{\star}}\rangle)^{2}.$

### v3

@@ 2.2 Separation between inductive and transductive settings
Proposition 7.
Let $d=1$. Consider the Vovk-Azoury-Warmuth algorithm, which computes, upon observing $x_{t}$ at time $t$,
@@
Proof of Proposition 7 We consider two cases. In the first case, the regularization is given at a large scale $\lambda\geq T$. For this case, consider $x_{t}=1$ and $y_{t}=1$ for all $t$. Then we have
@@
$=\left\lvert 1+\frac{T^{2}2^{t+1}\left((-1)^{t}+2^{t-1}\right)/3}{\lambda+4T^{2}(4^{t}-1)/3}\right\rvert$
@@
Although our result does not resolve the question in (Gaillard et al., 2019, Section 4.1), it hints, in view of Theorem 1, at a separation between the transductive setup and the inductive setup. In the following section, we show that the separation indeed exists for the clipped linear function class.
@@
$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\inf_{\theta\in\mathbb{R}}\left(\sum_{t=1}^{T}(y_{t}-\operatorname{clip_{m}}\left(x_{t}\theta\right))^{2}\right)=\Omega(T).$
Proof of Proposition 8 With a slight abuse of the notation, denote $x^{(i)}=10^{i}$ and $\theta^{(i)}=\frac{15}{16x^{(i)}}$ for $i\in\{0,1,\ldots,2^{T}-2\}$. We have
$\begin{cases}\theta^{(i)}x^{(j)}>7/8&\text{if\leavevmode\nobreak\ }j\geq i,\\ \theta^{(i)}x^{(j)}<1/8&\text{if\leavevmode\nobreak\ }j<i.\end{cases}$
@@
$\begin{split}\begin{cases}x_{t}\widetilde{\theta}>y_{t}=7/8&\text{if\leavevmode\nobreak\ }\hat{y}_{t}\leq 1/2,\\ x_{t}\widetilde{\theta}<y_{t}=1/8&\text{if\leavevmode\nobreak\ }\hat{y}_{t}>1/2\end{cases}\end{split}$ (17)
with the following binary searching construction. To start with, we set the left and the right end to be respectively $a_{1}=0$ and $b_{1}=2^{T}-2$. At each round $t$, the adversary presents the middle point of $a_{t}$ and $b_{t}$, that is, $i_{t}=\lfloor\frac{a_{t}+b_{t}}{2}\rfloor$ and $x_{t}=x^{(i_{t})}$. Then if $\hat{y}_{t}\leq 1/2$, then we set $y_{t}=7/8$, $a_{t+1}=i_{t}+1$, and $b_{t+1}=b_{t}$, else we set $y_{t}=1/8$, $a_{t+1}=a_{t}$, and $b_{t+1}=i_{t}-1$. At the beginning, we have $n_{1}=2^{T}-1$ covariates in between $a_{1}$ and $b_{1}$ and each time we are left with $(n_{t}-1)/2$ covariates in between $a_{t+1}$ and $b_{t+1}$. Then by induction, $n_{t}=2^{T-t+1}-1$. Thus, $n_{T}=1$. Finally, it is clear that $\widetilde{\theta}=\theta^{(i_{T})}$ satisfies (17).
@@ 3 Logistic Regression
For the transductive setting of log loss, it is well-known that the exact minimax risk is characterized by the Shtarkov sum (Shtarkov, 1987; Wu et al., 2022), and this exact minimax risk is achieved by the Normalized Maximum Likelihood (NML) algorithm. However, the Shtarkov sum is difficult to interpret, and the NML algorithm is computationally prohibitive, with the computational complexity scaling as $\Omega(2^{T})$ in the case of logistic regression. In terms of estimating the Shtarkov sum, Jacquet et al. (2022) notably proved that for the logistic loss, the Shtarkov sum scales as $\frac{d}{2}\log\frac{T}{2\pi}+\log\int\sqrt{\det(J_{F}(\theta;x_{1},\dots,x_{T}))}\,d\theta+o(1)$ when the covariates $x_{1},\dots,x_{T}$ are selected from a finite set of size $N=o(\sqrt{T})$ and dimension $d\leq N$, where $J_{F}(\theta;x_{1},\dots,x_{T})$ represents the Jeffreys prior given $x_{1},\dots,x_{T}$. This result leaves the interpretation of the Jeffreys prior term open for general covariates $x_{1},\dots,x_{T}$. For the special case where the covariates are uniformly sampled from the unit sphere, an asymptotic upper bound of $O(d\log T)$ for the Shtarkov sum can be inferred. Recent work by Drmota, Jacquet, Wu, and Szpankowski (2024b) claims to achieve an upper bound of $O(d\log(T/d))$ through a covering argument, which we formally detail in Appendix D but with worse constants in the regret bound. However, the covering-based argument lacks implementable algorithms for constructing the covering.
@@
We provide the first non-asymptotic, tight upper bound on the regret for transductive logistic regression without any assumptions on the covariates $x_{1},\dots,x_{T}$ or the optimal comparator $\theta^{\star}$}. Importantly, in Section 5, we discuss practical implementable versions of our algorithms that are guaranteed to converge to the optimal solution, with the only caveat that the number of iterations required may grow exponentially with the dimension in the worst-case scenario. In the specific case of random design logistic regression, we achieve in Section 6.2 a truly polynomial time implementation.
@@ 3.1 Background for exponential weights algorithm for logistic loss
Similar to the mixability of the squared loss (Lemma 6), the mixability of logistic loss facilitates the analysis of exponential weights algorithms (Vovk, 1990), or equivalently, Bayesian learning algorithms (Kakade and Ng, 2004). Formally, the logistic loss defined as $\ell_{\theta}(x,y)=-\log(\sigma(y\langle x,\theta\rangle))$ is $\eta$-mixable with $\eta=1$, i.e., the improper prediction of $\mathbb{E}_{\theta\sim\rho}\left[\sigma(y\langle x,\theta\rangle)\right]$ satisfies
$-\log\left(\mathbb{E}_{\theta\sim\rho}\left[\sigma(y\langle x,\theta\rangle)\right]\right)=-\log\left(\mathbb{E}_{\theta\sim\rho}\exp(-\ell_{\theta}(x,y))\right).$ (18)
@@
$\begin{split}&-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\nolimits_{t=1}^{T}\eta\ell_{\theta}(x_{t},y_{t})\right)\right)\\ &\approx-\frac{1}{\eta}\log\int\mu({\theta^{\star}})\exp\left(-\sum\limits_{t=1}^{T}\eta\ell_{{\theta^{\star}}}(x_{t},y_{t})-\frac{1}{2}(\theta-{\theta^{\star}})^{\top}\mathsf{H}_{t}({\theta^{\star}})(\theta-{\theta^{\star}})\right)d\theta,\end{split}$ (19)
where ${\theta^{\star}}\vcentcolon=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum\limits_{t=1}^{T}\ell_{\theta}(x_{t},y_{t})$ and $\mathsf{H}_{t}({\theta^{\star}})=\frac{x_{t}x_{t}^{\top}}{(1+\exp(({\theta^{\star}})^{\top}x_{t}))(1+\exp(-({\theta^{\star}})^{\top}x_{t}))}$ is the Hessian matrix of $\ell_{\theta}(x_{t},y_{t})$ at $\theta^{\star}$}. Applying an exact Laplace formula for quadratic forms (Lemma 16 below) to (LABEL:eq:laplace-approximation) together with Lemma 4 and Eq. 18, the asymptotic logistic loss of an exponential weights algorithm with prior $\mu$ is approximately
@@
Instead, Kakade and Ng (2004) employs a Gaussian prior and obtains an upper bound through quadratic approximation of the logistic loss. Concretely, by choosing the prior distribution $\mu=\mathcal{N}(0,\nu^{2}I)$ to be gaussian for some $\nu>0$ to be specified, by Lemmas 4, 5 and 18, the loss of the exponential weights algorithm satisfies
$\sum\limits_{t=1}^{T}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\left[\sigma(y_{t}\langle x_{t},\theta\rangle)\right]\right)=\inf_{\varphi}\left(-\sum\limits_{t=1}^{T}\mathbb{E}_{\theta\sim\varphi}\left[\log(\sigma(y_{t}\langle x_{t},\theta\rangle))\right]+\mathcal{KL}(\varphi\|\mu)\right).$
By choosing $\varphi=\mathcal{N}({\theta^{\star}},\varepsilon^{2}I)$ for some $\varepsilon>0$, together with $|\partial^{2}(-\log\sigma(y_{t}z))/\partial z^{2}|\leq 1/4$, the quadratic approximation of $\mathbb{E}_{\theta\sim\varphi}\sum\limits_{t=1}^{T}\log\sigma(y_{t}\langle x_{t},\theta\rangle)$ near $\theta^{\star}$} is upper bounded by
@@
In fact, $\theta^{\star}$} can be unbounded in the logistic regression case. Moreover, even $\min_{t}|\langle{\theta^{\star}},x_{t}\rangle|$ can be unbounded, as can be seen from the realizable case where all $x_{t}$’s have the same label. Indeed, since a lower bound for the inductive case is established by Foster et al. (2018, Theorem 5) in the following theorem, it is not possible to obtain our result with any prior that does not depend on $x_{1},\ldots,x_{T}$.
Theorem 9 (Foster et al. (2018)).
@@
This theorem, in conjunction with our upper bound Theorem 13, establishes a learnability separation in terms of assumptions between the transductive and inductive settings for logistic loss. Specifically, it demonstrates that the transductive setting is crucial for achieving assumption-free regret bounds. We note that Dzhamtyrova and Kalnishkan (2021) obtained an asymptotically tight lower bound for a related multiclass softmax problem under KL loss with the norm in the bound; while their setting differs, it would be interesting to see if similar ideas improve Theorem 9.
@@ 3.2 Slab-experts
For this comparator [^2] The parameter $\theta^{\star}$} is not always well-defined in $\mathbb{R}^{d}$. For instance, when the sequence is realizable, $\theta^{\star}$} will tend to be infinite in the separating direction. Since this is a minor technical issue, we avoid the complication in this paper. For a more rigorous treatment, we refer to Mourtada and Ga¨ıffas (2022). and any $1\leq t\leq T$, if $y_{t}\langle x_{t},{\theta^{\star}}\rangle\leq 0$, then $|\langle{\theta^{\star}},x_{t}\rangle|\leq T\log(2)$.
Proof of Lemma 10 For any $t$,
$\log(\sigma(y_{t}\langle x_{t},{\theta^{\star}}\rangle))\leq\sum_{t=1}^{T}-\log(\sigma(y_{t}\langle x_{t},{\theta^{\star}}{}\rangle))\leq\sum_{t=1}^{T}-\log(\sigma(0))=T\log 2,$
@@
Lemma 10 states that for any $x_{t}$ wrongly predicted covariate by $\theta^{\star}$}, the margin $|\langle{\theta^{\star}},x_{t}\rangle|$ will not exceed $T$. This motivates the following separation of the covariates: either $|\langle{\theta^{\star}},x_{t}\rangle|\leq T$ or $y_{t}\langle{\theta^{\star}},x_{t}\rangle>T$.
In the following, we first introduce a partial exponential weights algorithm that depends on any chosen subsets of the covariates that only predicts on the chosen subsets. Then, we use these partial exponential weights algorithms to build slab-experts.
Definition 11 (Partial exponential weights).
@@
Definition 12 (Slab-expert).
A slab-expert $B_{\theta,\alpha}$ is an expert that is specified by a vector $\theta\in\mathbb{R}^{d}$ and a regularization parameter $\alpha>0$. Given $x_{1},\ldots,x_{T}$, the expert constructs a slab $K_{\theta}=\{x_{t}:\left\lvert\langle x_{t},\theta\rangle\right\rvert\leq T\}$ and predicts labels in the following way: At time step $t$, if $x_{t}\in K_{\theta}$, the expert predicts on $x_{t}$ according to $\mathrm{EWA}_{K_{\theta},\alpha}$ with logistic loss and learning rate $1$. If $x_{t}\notin K_{\theta}$, then the expert predicts $\hat{y}_{t}=\operatorname{sign}(\langle x_{t},\theta\rangle)$ with probability $1$. We denote by $B_{\theta,\alpha}(x_{t},y_{t})$ the probability of predicting $y_{t}$ upon observing $x_{t}$.
@@
Construct the set of all unique slab-experts $\mathcal{A}=\{B_{\theta,\alpha}\mid{}\theta\in\mathbb{R}^{d}\}$ (removing copies) with logistic loss and learning rate $1$.
@@
   The learner predicts with aggregation on the slab-experts $\tilde{\rho}_{t}(x_{t},\cdot)=\sum\limits_{B\in\mathcal{A}}p_{t}(B)B(x_{t},\cdot)\text{\leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ }p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}\log B(x_{i},y_{i})\right)$ (22) and $B(x_{i},y_{i})$ is probability of $B$ predicting $y_{i}$ on $x_{i}$.
@@
$-\sum\limits_{t=1}^{T}\log\tilde{\rho}_{t}(x_{t},y_{t})\leq-\sum\limits_{t=1}^{T}\log(\sigma(y_{t}\langle x_{t},\theta^{\star}\rangle))+6d\log(eT).$
There are several noteworthy aspects of this result. First and foremost, it provides the first assumption-free guarantee, meaning that no assumptions are required on the covariates $x_{1},\dots,x_{T}$ or the comparator $\theta^{\star}$}. In Appendix D, we will present an alternative covering approach that achieves the same upper bound, where the covering is constructed in a manner analogous to the slab division on the covariates. However, our focus remains on the slab-experts approach, as the covering approach lacks implementable algorithms for constructing the covering. Indeed, our algorithm can be implemented in time polynomial in $T$, though exponential in $d$ in the worst case. This efficiency stems from the slab-experts themselves, each of which consists of an exponential weight algorithm with a Gaussian prior and a deterministic linear separator. The exponential weight algorithm with a Gaussian prior can be computed in polynomial time through log-concave sampling, and the aggregation of the slab-experts can be performed in $\mathrm{poly}(T^{d})$ time. Importantly, under mild additional assumptions on the distribution in the i.i.d. batch setup, we can reduce the number of experts to $2$ with pretraining, providing a polynomial time algorithm. For further details, see Sections 5 and 6. Additionally, the slab-experts approach has two notable aspects:
@@
Lemma 14.
If $T\geq d$, then there are at most $(eT/d)^{4d}$ distinct slab-experts.
Proof of Lemma 14 By the Vapnik-Chervonenkis-Sauer-Shelah lemma (Shalev-Shwartz and Ben-David (2014, Lemma 6.10)), the number of all possible linear separations on the set $K$ is upper bounded by $(eT/(d+1))^{(d+1)}\leq(eT/d)^{2d}$, whenever $T\geq d$. Therefore, since slab-experts can be formulated as intersections of two linear separations on the set $K$, the number of slab-experts is upper bounded by $(eT/d)^{4d}$. ∎
@@
Proposition 15.
Let $K\subset\{x_{t}\}_{t\in[T]}$ and ${\theta^{\star}}\in\mathbb{R}^{d}$. Then the partial exponential weights algorithm $\mathrm{EWA}_{K,\alpha}$ as defined in Definition 11 with prediction restricted on covariates in $K$ satisfies a loss upper bound of
$\hskip-72.26999pt-\sum_{x\in K}\log\left(\mathbb{E}_{\theta\sim\rho_{x}}\left[\sigma(y_{x}\langle x,\theta\rangle)\right]\right)$
$\leq-\sum_{x\in K}\log(\sigma(y_{x}\langle x,\theta^{\star}\rangle))+\alpha({\theta^{\star}})^{\top}\left(\sum\limits_{x\in K}xx^{\top}\right){\theta^{\star}}+\frac{d}{2}\log\left(1+\frac{1}{8\alpha}\right)$
on the set $K$ where $\rho_{x}$ is the prediction of $\mathrm{EWA}_{K,\alpha}$ on $x$. Specifically, suppose that $\theta^{\star}$} as defined in Eq. 21 satisfies $|\langle x,{\theta^{\star}}\rangle|\leq T$ for all $x\in K$ and let $\alpha=1/T^{3}$. We have
@@
Lemma 16 (Exact Laplace formula for quadratic forms).
Let $Q(\theta)=\theta^{\top}A\theta+b^{\top}\theta+c$, where $A\in\mathbb{R}^{d\times d}$ is a positive definite matrix, $b\in\mathbb{R}^{d}$ and $c\in\mathbb{R}$. Then
$\int\limits_{\mathbb{R}^{d}}\exp(-Q(\theta))d\theta=\exp\left(-\inf\limits_{\theta\in\mathbb{R}^{d}}Q(\theta)\right)\frac{\pi^{d/2}}{\sqrt{\det(A)}}.$
Proof of Proposition 15 By Lemma 4, we have
$-\sum_{x\in K}\log\left(\mathbb{E}_{\theta\sim\rho_{x}}\left[\sigma(y_{x}\langle x,\theta\rangle)\right]\right)=-\log\left(\int\limits_{\mathbb{R}^{d}}\exp\left(\sum\limits_{x\in K}\log(\sigma(y_{x}\langle x,\theta\rangle))\right)\mu_{K,\alpha}(d\theta)\right).$
Let $f(z)=-\log(\sigma(z))$. Since $|f^{\prime\prime}(z)|\leq 1/4$, using the Taylor expansion for $f$ we have for any $x\in K$,
$\log(\sigma(y_{x}\langle x,\theta\rangle))\leq-\log(\sigma(y_{x}\langle x,{\theta^{\star}}\rangle))-\frac{1}{1+\exp(y_{x}\langle x,{\theta^{\star}}\rangle)}\cdot(y_{x}\langle x,\theta-{\theta^{\star}}\rangle)+\frac{1}{8}(\langle x,\theta-{\theta^{\star}}\rangle)^{2}.$